OpenStudy (anonymous):

The line y = mx + m intersects the graph of y = x^2 at x = m + 2. Find the value of m.

OpenStudy (anonymous):

i guess this means \[m(m+2)+m=(m+2)^2\] so we can solve that for \(m\)

OpenStudy (anonymous):

\[m^2+3m=m^2+4m+2\] \[3m=4m+2\] \[m=-2\]

OpenStudy (anonymous):

AWESOME! Thank you soooo much!

OpenStudy (anonymous):

yw, hope it is clear why it is true

OpenStudy (anonymous):

Yes, I see it now. :)

OpenStudy (anonymous):

u mean -4

OpenStudy (anonymous):

Yes thank you